OpenStudy (anonymous):

A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $120 for 100 minutes of usage and $95 for 50 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x? Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 0.50 Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 2 Draw a graph which joins the points (120, 100) and (95, 50)

OpenStudy (anonymous):

Well, we are told that the x stands for total minutes used and y for the charge. We know that x=100 means that y=120. So the line must pass through (100, 120). We know that x=50 means y=95. So it should also pass through (50, 95). Now we have to calculate the slope. To do so we basically have to see the ratio between the difference in the y and the difference of the x of two points. Say we denote our points as A and B: $$ A = (100, 120) \\ B = (50, 90) \\ \Delta x = A_x - B_x = 100 - 50 = 50\\ \Delta y = A_y - B_y = 120 - 95 = 25\\ \text{slope} = \frac{\Delta y}{\Delta x} = \frac{25}{50} = \frac{1}{2} $$So we have a slope of 0.5